.What is the magnification when an object is placed at 2f of a convex mirror a) 1/3 b) 2/3 c) 1 d) 3/2

Grade:upto college level

2 Answers

Saurabh Koranglekar
3 years ago
According to the problem the distance from mirror to object is 2f

and the focal length is f

Now, we know that

1/v+ 1/u =1/f [ v is the distance of image from mirror, u is the distance of object from the mirror and f is the focal length]

1/v+ 1/(-2f) =1/f

1/v=3/2f

v=2f/3

we know that magnification can be calculated:

height of image/ height of object

magnification, m=2f/3/2f

=1/3

Vikas TU
3 years ago
U=-2f
V=?
1/u+1/v=1/f
1/-2f + 1/v = 1 /f
1/v= 1/f + 1/2f
V = 2f/3
Now, m=-v/u
=-2f/3*-2f
=1/3................................